OpenStudy (anonymous):

Two vehicles collide and stick together. After the collision, their combined y-momentum is 2.40 x 10^4 kilogram meters/second, and their x-momentum is 7.00 x 10^4 kilogram meters/second. What is the angle of the motion of the two vehicles, with respect to the x-axis? Could someone please assist? EDIT: Okay, so I tried myself and here's what I did: 2.40 x 10^4 * 7.00 x 10^4 = 168 = 16.8 degrees Not sure if right, anyone want to confirm?

OpenStudy (astrophysics):

I'm not sure what you're doing, but momentum is \[\vec p = m \vec v\] you should make a drawing for initial case and final case, are you given directions?

OpenStudy (astrophysics):

Anyways to find the angle of the motion we use \[\tan \theta = \left( \frac{ |P_y f| }{ |P_x f| } \right)\]

OpenStudy (anonymous):

I wasn't given any specific directions, all I wrote was all that was there... Ah, alright, so it'll be tan(theta) = (2.40 x 10^4 (f) / 7.00 x 10^4 (f))?

OpenStudy (astrophysics):

The f here just means final case, but yes, that's fine \[\tan \theta = \left( \frac{ 2.4 }{ 7 } \right)\]

OpenStudy (astrophysics):

Do you know how to isolate angle theta?

OpenStudy (astrophysics):

By the way the reason I used 2.4 and 7 and did not include 10^4 because 10^4/10^4 is just 1.

OpenStudy (anonymous):

theta^-1 right?

OpenStudy (astrophysics):

Right, we take the inverse! :)

OpenStudy (anonymous):

Oh wait...would it look something like this? (theta) = tan^-1= (2.4/7)

OpenStudy (astrophysics):

\[\theta = \tan ^{-1} \left( \frac{ 2.4 }{ 7 } \right)\]

OpenStudy (astrophysics):